Circular No. 8548

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2005cm IN IC 5142 Further to IAUC 8541, L. A. G. Monard reports his discovery of an ap-

parent supernova on unfiltered images taken on June 20.124 (at mag ∼ 16.9) and 20.779 UT (mag ∼ 16.8). SN 2005cm is located at α = 21^{h}55^{m}17^{s}.72, δ = −65^{o}30′48′′.3 (equinox 2000.0), which is 16′′ west and 13′′ south of the nucleus of IC 5142. Nothing is visible at this location on an image taken by Monard on June 5.085 (limiting red mag 18.5) or on an image of the Digitized Sky Survey (limiting red mag 20.5).

COMETS C/2005 H4–H9 (SOHO) Following are the initial available positions for additional comets found

on SOHO website images (continuation to IAUC 8547); all are Kreutz sungrazers except for C/2005 H7 (no known group membership) and C/2005 H9 (Meyer group). K. Battams notes that C/2005 H4 was faint and diffuse (yielding only five images due to data gaps), peaking in brightness at mag ∼ 8.1 at 6.4R_{⊙} (Apr. 21.993 UT). C/2005 H5 had a diffuse ‘teardrop’ shape, reaching mag ∼ 7.2 at 6.8R_{⊙} (Apr. 24.851). C/2005 H6 was very faint and elongated; it reached mag ∼ 8.1 at 5.5R_{⊙} (Apr. 25.129). C/2005 H7 was very small (one or two pixels in size) and stellar, reaching mag ∼ 7.4 at 6.2R_{⊙} (Apr. 25.226). C/2005 H8 was very small and diffuse, reaching mag ∼ 7.9 at 6.2R_{⊙} (Apr. 26.268). C/2005 H9 was diffuse and elongated, measuring ∼ 34′′ in length when it reached peak brightness at mag ∼ 7.4 (at 7.0R_{⊙}) on Apr. 29.018.

Comet 2005 UT α_{2000} δ_{2000} Inst. F MPEC
C/2005 H4 Apr. 21.976 2^{h}05.^{m}8 +11^{o}13′ C2 CL 2005-K44
C/2005 H5  24.826 2 17.1 +12 07 C2 TH 2005-K44
C/2005 H6  25.101 2 14.7 +11 51 C2 RK 2005-K44
C/2005 H7  25.201 2 04.8 +13 00 C2 RK 2005-K44
C/2005 H8  26.226 2 21.7 +12 28 C2 BZ 2005-K44
C/2005 H9  28.811 2 30.3 +15 26 C2 BZ 2005-M14

COMET C/2005 K2 (LINEAR) J. A. Farrell, Jamez Springs, NM, reports the following separations

(and 1σ errors), all in p.a. 47^{o}.4, of the secondary component to the main nuclear condensation of this comet (cf. IAUC 8543, 8545) from CCD images taken with a 0.4-m reflector: June 13.17 UT, 38′′.1 ± 0′′.7; 14.17, 40′′.0 ± 0′′.4; 17.16, 43′′.1 ± 0′′.8.

2005 June 20  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2005 CBAT  Daniel W. E. Green